# On the cohomological dimension of Siegel modular varieties and the modularity of formal Siegel modular forms

### v1

@@ 5 Main results
In this section, we first show that the boundary of the minimal compactification of Siegel modular variety satisfies the weak G3 properties with respect to the Hodge line bundle $\omega$, and then apply the results in the previous section to show our main theorems.
@@
Bruinier and Raum defined the notion of formal Siegel modular forms of weight $k$ and cogenus $l$ in Section $3$ in [BR24]. The complex vector space of formal Siegel modular forms of weight $k$, cogenus $l$ and level $\Gamma$ is denoted as $\mathrm{FM}^{(g,l)}_{k}(\Gamma)$. We shall not recall their definition here, since they have proved that $\mathrm{FM}_{k}^{(g,1)}(\Gamma)$ can be realized as the global section of $\omega^{k}$ on the formal completion of $A_{g,\Gamma}^{\min}$ along the boundry. More precisely, let $\hat{A}^{\min}_{g,\Gamma}=(A^{\min}_{g,\Gamma})_{/D^{\min}_{\Gamma}}$ denote the formal completion and $\hat{\omega}$ be the pullback of $\omega$ on $\hat{A}^{\min}_{g,\Gamma}$.
@@
We may give a proof of the dimension bound of formal sections by using algebraic geometric language in a forthcoming paper (joint with Peihang Wu, to deal with general cases), as we have certain analogues of the results about the existence of slope bound and theta decomposition. However, in this paper, we simply apply the result on the dimension bound of formal Siegel modular forms of level $1$.
@@
Since $A_{g}^{\min}$ is normal, we have $\mathcal{O}_{\hat{A}_{g}^{\min}}$ is integrally closed in $\mathcal{M}_{\hat{A}_{g}^{\min}}$ hence $\mathcal{O}_{\hat{A}_{g}^{\min}}[T]$ is integrally closed in $\mathcal{M}_{\hat{A}_{g}^{\min}}[T]$. As $\bigoplus_{m\geq 0}\hat{\omega}^{m}$ is locally isomorphic to $\mathcal{O}_{\hat{A}_{g}^{\min}}[T]$, we also have $\bigoplus_{m\geq 0}\hat{\omega}^{m}$ is integrally closed in $\mathcal{M}_{\hat{A}_{g}^{\min}}[T]$. Hence $A$ is integrally closed in $K(\hat{A}_{g}^{\min},\hat{\omega})[T]$. Since both quotient fields have same transcendence degree $d+1$, we must have
@@
By Proposition 3.6, it’s sufficient to show the claim in the minimal compactification cases. We first show that for any two arithmetic level groups $\Gamma_{1}\subset\Gamma_{2}$, the boundary $D^{\min}_{\Gamma_{1}}$ is weakly G2 with respect to $\omega$ in $A_{g,\Gamma_{1}}^{\min}$ if and only if $D^{\min}_{\Gamma_{2}}$ is weakly G2 with respect to $\omega$ in $A_{g,\Gamma_{2}}^{\min}$. Choose an arithmetic subgroup $\Gamma_{0}\subset\Gamma_{1}\subset\Gamma_{2}$ that is normal in $\Gamma_{2}$. Then we have $A_{g,\Gamma_{i}}^{\min}=(A_{g,\Gamma_{0}}^{\min})/G_{i}$ where $G_{i}=\Gamma_{i}/\Gamma_{0}$, and the canonical maps between the minimal compactifications are exactly the natural quotient maps.
@@
Applying Proposition 3.5, we have $D^{\min}_{\Gamma_{1}}$ is weakly G2 with respect to $\omega$ in $A_{g,\Gamma_{1}}^{\min}$ if and only if $D^{\min}_{\Gamma_{0}}$ is weakly G2 with respect to $\omega$ in $A_{g,\Gamma_{0}}^{\min}$, which is further equivalent to $D^{\min}_{\Gamma_{2}}$ is weakly G2 with respect to $\omega$ in $A_{g,\Gamma_{2}}^{\min}$. The claim of this corollary follows from that $\Gamma\cap\Gamma(1)$ is an arithmetic group.
@@
By the main theorem of [BMS67], $\Gamma_{0}$ contains a principal congruence subgroup $\Gamma(N)$ for some $N$, and is therefore an arithmetic level subgroup itself. Under the Galois correspondence, $\Gamma_{0}$ corresponds to the Siegel modular variety $A_{g,\Gamma_{0}}$, and hence $B\cong A_{g,\Gamma_{0}}$. ∎
@@
Without loss of generality, we can further assume $\Gamma_{1}$ is normal in $\Gamma_{2}$, and we denote $G=\Gamma_{2}/\Gamma_{1}$ which acts freely on $A_{g,\Gamma_{1}}$. Applying Theorem 3.7, we have Cartesian square
@@
By [Sta25, Tag 0476] and the flatness of $\pi_{\Gamma_{1},\Gamma_{2}}$, we have $\tilde{\pi}_{\Gamma_{1},\Gamma_{2}}^{-1}(B_{\Gamma_{2}}\backslash\Delta_{\Gamma_{2}}^{\prime})=B_{\Gamma_{1}}\backslash\Delta_{\Gamma_{1}}^{\prime}$, hence we have $\pi_{\Gamma_{1},\Gamma_{2}}^{-1}(\Delta_{\Gamma_{2}})=\Delta_{\Gamma_{1}}$.
Now we can conclude that $\Delta_{\Gamma}$ is empty because the Hecke orbit of every closed point $z$ in $A_{g,\Gamma}$ is Zariski dense but $\Delta_{\Gamma}$ is contained in a codimension $1$ closed subvariety. Therefore $f_{\Gamma}:B_{\Gamma}\to A_{g,\Gamma}$ is finite étale, hence it coincides $\pi_{\Gamma_{0},\Gamma}:A_{g,\Gamma_{0}}\to A_{g,\Gamma}$ for some $\Gamma_{0}\subset\Gamma$ by Proposition 5.5. By definition $B_{\Gamma}^{\min}$ is the relative normalization of $A^{\min}_{g,\Gamma}$ in $K(B_{\Gamma})=K(A_{g,\Gamma_{0}})$, thus $B_{\Gamma}^{\min}=A_{g,\Gamma_{0}}^{\min}$. So we have $B_{\Gamma}^{\min}\backslash B_{\Gamma}$ is connected thus it has to be $i^{\prime}(D_{\Gamma}^{\min})$. As $f_{\Gamma}:B^{\min}_{\Gamma}\to A_{g,\Gamma}^{\min}$ is étale in $B_{\Gamma}$ and an open neighborhood of $B_{\Gamma}^{\min}\backslash B_{\Gamma}$, it is étale on the whole variety $B_{\Gamma}^{\min}$. Since $f_{\Gamma}$ is an isomorphism on $f^{-1}_{\Gamma}(D_{\Gamma}^{\min})$, we have $f_{\Gamma}$ is an isomorphism hence $D_{\Gamma}^{\min}$ is weakly G3 in $A_{g,\Gamma}^{\min}$ with respect to $\omega$. ∎
@@
By Theorem 5.7, every effective divisor intersects nontrivially with the boundary, thus the condition $\mathrm{Lef}(A_{g,\Gamma}^{\min},D^{\min}_{\Gamma})$ follows from Proposition 4.8 and Theorem 5.6. For the pair $(A_{g,\Gamma}^{\Sigma},D^{\Sigma}_{\Gamma})$, we need the weak G3 property with respect to the ample line bundle $\omega^{k}(-D^{\Sigma}_{\Gamma})$ where $k$ is an integer large enough. This follows from the composition of the inculsion maps
@@
For general cases, choose a small enough normal arithmetic subgroup $\Gamma^{\prime}\subset\Gamma$, we have proved $\operatorname{ccd}(A_{g,\Gamma^{\prime}})\leq d-2$. Denote $G=\Gamma/\Gamma^{\prime}$, we have $A_{g,\Gamma}$ is the coarse moduli of the tame quotient stack $A_{g,\Gamma^{\prime}}/G]$, thus we have the spectral sequence
@@
where $i:[A_{g,\Gamma^{\prime}}/G]\to A_{g,\Gamma}$ is the map from a stack to its coarse moduli, and the last equality is given by $i_{*}$ is exact and $i_{*}i^{*}\omega_{\Gamma}^{k}=\omega_{\Gamma}^{k}\otimes i_{*}\mathcal{O}_{[A_{g,\Gamma^{\prime}}/G]}=\omega_{\Gamma}^{k}$. Since $G$ is finite, the classical result is that $H^{p}(G,M)$ is $|G|$-torsion for $p>0$, hence vanish if $M$ is uniquely $|G|$-divisible. So we have $E^{p,q}_{2}=0$ if $p>0$, thus
@@
restricted from $\pi$ has relative dimension dl(l+1)/21d-l(l+1)/2-1. Apply Theorem 5.10 we have
@@
where $\mathcal{H}^{s}_{F_{\Gamma,j}^{\Sigma}}(A_{g,\Gamma}^{\Sigma}\backslash D_{\Gamma,j-1}^{\Sigma},\mathcal{L})$ is the sheaf of cohomology with support, see [Sta25, Tag 0A39] for precise definition. As $D_{\Gamma}^{\Sigma}$ is a Cartier divisor, we have $\mathcal{H}^{s}_{F_{\Gamma,j}^{\Sigma}}(A_{g,\Gamma}^{\Sigma}\backslash D_{\Gamma,j-1}^{\Sigma},\mathcal{L})=0$ unless $s=1$. Since $\operatorname{ccd}(F_{\Gamma,j}^{\Sigma})\leq g(g+1)/2-3$, we have $E_{2}^{r,s}=0$ if r>g(g+1)/23r>g(g+1)/2-3 and $s\neq 1$. Thus in the first spectral sequence, $E_{1}^{p,q}=0$ if p+q>d2p+q>d-2. Therefore, $H^{i}(A_{g,\Gamma}^{\Sigma}\backslash D_{\Gamma,l}^{\Sigma},\mathcal{L})=0$ if i>d2i>d-2, hence $\operatorname{ccd}(A_{g,\Gamma}^{\Sigma}\backslash D_{\Gamma,l}^{\Sigma})\leq d-2$ by Lemma 4.2. ∎
@@
Choose a normal arithmetic subgroup $\Gamma^{\prime}\subset\Gamma$ small enough such that there exists a smooth toroidal compactification $A_{g,\Gamma^{\prime}}^{\Sigma}$. Then by the previous proposition and Proposition 4.5, we have that every effective divisor in $A_{g,\Gamma^{\prime}}^{\Sigma}$ intersects nontrivially with $D^{\Sigma}_{\Gamma^{\prime},l}$ and $\mathrm{Lef}(A_{g,\Gamma^{\prime}}^{\Sigma},D^{\Sigma}_{\Gamma^{\prime},l})$. Thus every effective divisor in $A_{g,\Gamma}^{\min}$ intersects nontrivially with $D^{\min}_{\Gamma,l}$, and by Proposition 4.8, $D^{\Sigma}_{\Gamma^{\prime},l}$ is weakly G3 in $A_{g,\Gamma^{\prime}}^{\Sigma}$ with respect to $\omega$. By Theorem of formal functions, we have
@@
The minimal compactification case follows from Corollary 5.12 and Proposition 4.7. For the toroidal compactification case, we have $D^{\Sigma}_{\Gamma,l}$ is weakly G3 in $A_{g,\Gamma}^{\Sigma}$ with respect to $\omega$ by Proposition 3.6, and then weakly G3 with respect to $\omega^{k}(-D^{\Sigma}_{\Gamma})$ by the same argument as in Theorem 5.9. Therefore $\mathrm{Lef}(A_{g,\Gamma}^{\Sigma},D^{\Sigma}_{\Gamma,l})$ follows from Proposition 4.7. ∎

### v2

@@ 5 Main results
In this section, we first show that the boundaries of the minimal compactifications of Siegel modular varieties satisfy the weak G3 properties with respect to the Hodge line bundle $\omega$, and then apply the results in the previous section to show our main theorems.
@@
Bruinier and Raum defined the notion of formal Siegel modular forms of weight $k$ and cogenus $l$ in Section $3$ in [BR24]. The complex vector space of the formal Siegel modular forms of weight $k$, cogenus $l$ and level $\Gamma$ is denoted as $\mathrm{FM}^{(g,l)}_{k}(\Gamma)$. We shall not recall their definition here, since they have proved that $\mathrm{FM}_{k}^{(g,1)}(\Gamma)$ can be realized as the global section of $\omega^{k}$ on the formal completion of $A_{g,\Gamma}^{\min}$ along the boundary. More precisely, let $\hat{A}^{\min}_{g,\Gamma}=(A^{\min}_{g,\Gamma})_{/D^{\min}_{\Gamma}}$ denote the formal completion and $\hat{\omega}$ be the pullback of $\omega$ on $\hat{A}^{\min}_{g,\Gamma}$.
@@
We hope to give a proof of the dimension bound of formal sections by using algebraic geometric language in a forthcoming paper (joint with Peihang Wu, to deal with general cases), as we have certain analogues of the results about the existence of slope bound and theta decomposition. However, in this paper, we simply apply the result on the dimension bound of formal Siegel modular forms of level $1$.
@@
Since $A_{g}^{\min}$ is normal, we see that $\mathcal{O}_{\hat{A}_{g}^{\min}}$ is integrally closed in $\mathcal{M}_{\hat{A}_{g}^{\min}}$ and hence $\mathcal{O}_{\hat{A}_{g}^{\min}}[T]$ is integrally closed in $\mathcal{M}_{\hat{A}_{g}^{\min}}[T]$. As $\bigoplus_{m\geq 0}\hat{\omega}^{m}$ is locally isomorphic to $\mathcal{O}_{\hat{A}_{g}^{\min}}[T]$, we also see that $\bigoplus_{m\geq 0}\hat{\omega}^{m}$ is integrally closed in $\mathcal{M}_{\hat{A}_{g}^{\min}}[T]$. Hence $A$ is integrally closed in $K(\hat{A}_{g}^{\min},\hat{\omega})[T]$. Since both quotient fields have the same transcendence degree $d+1$, we must have
@@
By Proposition 3.6, it suffices to show the claim in the minimal compactification cases. We first show that for any two arithmetic level groups $\Gamma_{1}\subset\Gamma_{2}$, the boundary $D^{\min}_{\Gamma_{1}}$ is weakly G2 with respect to $\omega$ in $A_{g,\Gamma_{1}}^{\min}$ if and only if $D^{\min}_{\Gamma_{2}}$ is weakly G2 with respect to $\omega$ in $A_{g,\Gamma_{2}}^{\min}$. Choose an arithmetic subgroup $\Gamma_{0}\subset\Gamma_{1}\subset\Gamma_{2}$ that is normal in $\Gamma_{2}$. Then we have $A_{g,\Gamma_{i}}^{\min}=(A_{g,\Gamma_{0}}^{\min})/G_{i}$ where $G_{i}=\Gamma_{i}/\Gamma_{0}$, and the canonical maps between the minimal compactifications are exactly the natural quotient maps.
@@
Applying Proposition 3.5, we see that $D^{\min}_{\Gamma_{1}}$ is weakly G2 with respect to $\omega$ in $A_{g,\Gamma_{1}}^{\min}$ if and only if $D^{\min}_{\Gamma_{0}}$ is weakly G2 with respect to $\omega$ in $A_{g,\Gamma_{0}}^{\min}$, which is further equivalent to $D^{\min}_{\Gamma_{2}}$ is weakly G2 with respect to $\omega$ in $A_{g,\Gamma_{2}}^{\min}$. The claim of this corollary follows from the fact that $\Gamma\cap\Gamma(1)$ is an arithmetic group.
@@
By the main theorem of [BMS67], $\Gamma_{0}$ contains a principal congruence subgroup $\Gamma(N)$ for some $N$ and is therefore an arithmetic level subgroup itself. Under the Galois correspondence, $\Gamma_{0}$ corresponds to the Siegel modular variety $A_{g,\Gamma_{0}}$, and hence $B\cong A_{g,\Gamma_{0}}$. ∎
@@
Without loss of generality, we can further assume $\Gamma_{1}$ is normal in $\Gamma_{2}$, and we denote $G=\Gamma_{2}/\Gamma_{1}$ which acts freely on $A_{g,\Gamma_{1}}$. Applying Theorem 3.7, we have a Cartesian square
@@
By [Sta25, Tag 0476] and the flatness of $\pi_{\Gamma_{1},\Gamma_{2}}$, we have $\tilde{\pi}_{\Gamma_{1},\Gamma_{2}}^{-1}(B_{\Gamma_{2}}\backslash\Delta_{\Gamma_{2}}^{\prime})=B_{\Gamma_{1}}\backslash\Delta_{\Gamma_{1}}^{\prime}$, and hence we have $\pi_{\Gamma_{1},\Gamma_{2}}^{-1}(\Delta_{\Gamma_{2}})=\Delta_{\Gamma_{1}}$.
Now we can conclude that $\Delta_{\Gamma}$ is empty because the Hecke orbit of every closed point $z$ in $A_{g,\Gamma}$ is Zariski dense but $\Delta_{\Gamma}$ is contained in a closed subvariety of codimension $1$. Therefore, $f_{\Gamma}:B_{\Gamma}\to A_{g,\Gamma}$ is finite étale and then coincides $\pi_{\Gamma_{0},\Gamma}:A_{g,\Gamma_{0}}\to A_{g,\Gamma}$ for some $\Gamma_{0}\subset\Gamma$ by Proposition 5.5. By definition $B_{\Gamma}^{\min}$ is the relative normalization of $A^{\min}_{g,\Gamma}$ in $K(B_{\Gamma})=K(A_{g,\Gamma_{0}})$, thus $B_{\Gamma}^{\min}=A_{g,\Gamma_{0}}^{\min}$. So we have $B_{\Gamma}^{\min}\backslash B_{\Gamma}$ is connected thus it has to be $i^{\prime}(D_{\Gamma}^{\min})$. As $f_{\Gamma}:B^{\min}_{\Gamma}\to A_{g,\Gamma}^{\min}$ is étale on $B_{\Gamma}$ and an open neighborhood of $B_{\Gamma}^{\min}\backslash B_{\Gamma}$, it is étale on the whole variety $B_{\Gamma}^{\min}$. Since $f_{\Gamma}$ is an isomorphism on $f^{-1}_{\Gamma}(D_{\Gamma}^{\min})$, we find that $f_{\Gamma}$ is an isomorphism and hence $D_{\Gamma}^{\min}$ is weakly G3 in $A_{g,\Gamma}^{\min}$ with respect to $\omega$. ∎
@@
By Theorem 5.7, every effective divisor intersects nontrivially with the boundary, thus the condition $\mathrm{Lef}(A_{g,\Gamma}^{\min},D^{\min}_{\Gamma})$ follows from Proposition 4.8 and Theorem 5.6. For the pair $(A_{g,\Gamma}^{\Sigma},D^{\Sigma}_{\Gamma})$, we need the weak G3 property with respect to the ample line bundle $\omega^{k}(-D^{\Sigma}_{\Gamma})$ where $k$ is an integer large enough. This follows from the fact that the composition of the inculsion maps
@@
For general cases, choose a small enough normal arithmetic subgroup $\Gamma^{\prime}\subset\Gamma$, we have proved $\operatorname{ccd}(A_{g,\Gamma^{\prime}})\leq d-2$. Denoting $G=\Gamma/\Gamma^{\prime}$, we have $A_{g,\Gamma}$ is the coarse moduli of the tame quotient stack $A_{g,\Gamma^{\prime}}/G]$, thus we have the spectral sequence
@@
where $i:[A_{g,\Gamma^{\prime}}/G]\to A_{g,\Gamma}$ is the map from a stack to its coarse moduli, and the last equality is given by the fact that $i_{*}$ is exact and $i_{*}i^{*}\omega_{\Gamma}^{k}=\omega_{\Gamma}^{k}\otimes i_{*}\mathcal{O}_{[A_{g,\Gamma^{\prime}}/G]}=\omega_{\Gamma}^{k}$. Since $G$ is finite, the classical result is that $H^{p}(G,M)$ is $|G|$-torsion for $p>0$, hence vanish if $M$ is uniquely $|G|$-divisible. So we have $E^{p,q}_{2}=0$ if $p>0$, thus
@@
restricted from $\pi$ has relative dimension dl(l+1)/21d-l(l+1)/2-1. Applying Theorem 5.10 we have
@@
where $\mathcal{H}^{s}_{F_{\Gamma,j}^{\Sigma}}(A_{g,\Gamma}^{\Sigma}\backslash D_{\Gamma,j-1}^{\Sigma},\mathcal{L})$ is the sheaf of cohomology with support, see [Sta25, Tag 0A39] for a precise definition. As $D_{\Gamma}^{\Sigma}$ is a Cartier divisor, we have $\mathcal{H}^{s}_{F_{\Gamma,j}^{\Sigma}}(A_{g,\Gamma}^{\Sigma}\backslash D_{\Gamma,j-1}^{\Sigma},\mathcal{L})=0$ unless $s=1$. Since $\operatorname{ccd}(F_{\Gamma,j}^{\Sigma})\leq g(g+1)/2-3$, we have $E_{2}^{r,s}=0$ if r>g(g+1)/23r>g(g+1)/2-3 and $s\neq 1$. Thus in the first spectral sequence, $E_{1}^{p,q}=0$ if p+q>d2p+q>d-2. Therefore, $H^{i}(A_{g,\Gamma}^{\Sigma}\backslash D_{\Gamma,l}^{\Sigma},\mathcal{L})=0$ if i>d2i>d-2, hence $\operatorname{ccd}(A_{g,\Gamma}^{\Sigma}\backslash D_{\Gamma,l}^{\Sigma})\leq d-2$ by Lemma 4.2. ∎
@@
Choose a normal arithmetic subgroup $\Gamma^{\prime}\subset\Gamma$ small enough such that there exists a smooth toroidal compactification $A_{g,\Gamma^{\prime}}^{\Sigma}$. Then by the previous proposition and Proposition 4.5, we have that every effective divisor in $A_{g,\Gamma^{\prime}}^{\Sigma}$ intersects nontrivially with $D^{\Sigma}_{\Gamma^{\prime},l}$ and $\mathrm{Lef}(A_{g,\Gamma^{\prime}}^{\Sigma},D^{\Sigma}_{\Gamma^{\prime},l})$. Thus, every effective divisor in $A_{g,\Gamma}^{\min}$ intersects nontrivially with $D^{\min}_{\Gamma,l}$, and by Proposition 4.8, $D^{\Sigma}_{\Gamma^{\prime},l}$ is weakly G3 in $A_{g,\Gamma^{\prime}}^{\Sigma}$ with respect to $\omega$. By Theorem of formal functions, we have
@@
The minimal compactification case follows from Corollary 5.12 and Proposition 4.7. For the toroidal compactification case, we have that $D^{\Sigma}_{\Gamma,l}$ is weakly G3 in $A_{g,\Gamma}^{\Sigma}$ with respect to $\omega$ by Proposition 3.6, and then weakly G3 with respect to $\omega^{k}(-D^{\Sigma}_{\Gamma})$ by the same argument as in Theorem 5.9. Therefore, $\mathrm{Lef}(A_{g,\Gamma}^{\Sigma},D^{\Sigma}_{\Gamma,l})$ follows from Proposition 4.7. ∎